# Typical Quantum States of the Universe are Observationally Indistinguishable

### v1

@@ 2.2 Main Motivation: The Past Hypothesis
In fact, it has been suggested for the explanation of the thermodynamic arrow of time (Feynman 1963, p. 115), (Penrose 1979, Lebowitz 1993, Albert 2000, Goldstein et al. 2020) that the initial state of the universe must be restricted to a set of low-entropy states; in a quantum theory, such a set would be given by a suitable subspace $\mathscr{H}_{0}$ of the Hilbert space $\mathscr{H}$ of the universe, and the presumed additional law might then be formulated as follows:
@@
Even though $\mathscr{H}_{0}$ has much lower dimension than $\mathscr{H}_{eq}$, $\mathscr{H}_{0}$ is still, like all subspaces representing macro-states, a high-dimensional subspace. Realistically, we can expect $d_{0}>>10^{10^{20}}$. With respect to the normalized uniform measure $u_{0}$, we expect that typical wave functions in $\mathbb{S}(\mathscr{H}_{0})$ will evolve in a way such that $S_{qB}$ increases in the medium and the long run, satisfying the Developmental Conjecture formulated by Goldstein et al. (2020), a version of the Second Law of Thermodynamics for a quantum universe.
Hence, if we choose (say) $\varepsilon=10^{-200}$, then for any $E_{z}$, typical individual wave functions in $\mathscr{H}_{0}$ will all lead to nearly the same probability distribution as $\rho_{0}$. In the philosophical literature, $\rho_{0}$ has been called the Wentaculus density matrix, corresponding to the initial state of the Wentaculus theory (Chen 2021, 2024a).
@@ 3.1 General Distinctions
There are different kinds of observational indistinguishability (OI) discussed in the philosophical literature. We draw some general distinctions regarding (1) OI and empirical equivalence, and (2) in-principle and in-practice OI.
Empirical equivalence is a relation holding among certain physical theories. Two theories may be regarded as empirically equivalent when they make the same predictions about all possible observations. For example, one can explore whether BM is empirically equivalent to OQM, GRW, or Everettian quantum mechanics by checking whether BM makes the same empirical predictions as the latter.
In contrast, OI is primarily a relation holding among models in the same theory. In the philosophical literature, there is a large literature about the OI of certain models of general relativity. Manchak (2009) showed that for almost every general relativistic spacetime, we can construct another one in such a way that no observation will distinguish between the two spacetimes.
In this paper, we focus on OI as it is applied to models of quantum theory. Here, we shall take an interpretation-neutral approach, and discuss how to apply the result in different interpretations in §3.2. Nevertheless, what we show may also be relevant to the issue of empirical equivalence among quantum theories.
@@
The second distinction is between in-principle OI and in-practice OI. People often focus on exact equivalence of predictions (or retrodictions) about outcomes of observations. This is done in the general relativistic case in terms of exact agreement (up to an isometry transformation) of past lightcones of observers, and in quantum foundations in terms of exact equality of theoretical distributions. In many cases, exact equivalence is sufficient but not necessary for OI. We never did and never will probe the exact microscopic detail of the past lightcone. In practice we will quickly exhaust our resources long before reaching the in-principle limits. Nevertheless, to prove sharp mathematical results, it is sometimes easier to focus on in-principle limits. Of course, in-principle OI implies in-practice OI.
Observation typicality concerns the stronger sense of OI—in principle and not just in practice. We show that in-principle OI follows from approximate equality of probabilistic predictions. The approximate equality of predictions is to be distinguished from the approximation of frequency to chance. For example in Bohmian mechanics, we justify the quantum equilibrium hypothesis using a law-of-large-number argument (Dürr et al. 1992), with frequency converging to chance: typically (with respect to the quantum equilibrium distribution), the empirical distribution converges to the theoretical distribution. The result establishes exact equality of the chance functions (theoretical distributions). While exact equality of probabilities is sufficient for in-principle OI, our argument below shows that it is not necessary.
According to a standard criterion (Nielsen and Chuang 2010, p.86), two quantum states are perfectly distinguishable by some experiment if and only if they are orthogonal. If we do not insist on perfect distinguishability, we may say that two quantum states are reliably distinguishable by some experiment if and only if they are (at least) approximately orthogonal.
We can specialize the criterion for OI to a particular observation $E_{z}$. For two quantum states $\Psi_{1}$ and $\Psi_{2}$, we say that they are perfectly distinguishable by observing the outcome $Z=z$ if and only if $\langle\Psi_{1}|\,E_{z}\,|\Psi_{1}\rangle=0$ and $\langle\Psi_{2}|\,E_{z}\,|\Psi_{2}\rangle=1$ or vice versa; let us assume the former is the case. If we know that the actual quantum state is either $\Psi_{1}$ or $\Psi_{2}$, the outcome $Z=z$ is a perfect indicator of $\Psi_{2}$. The indicator is perfect in the sense that, in 100% of the worlds compatible with the outcome $Z=z$, it correctly indicates the actual quantum state. If we do not insist on perfect distinguishability, we can allow the probabilities to be approximately zero or one, so that they are reliably distinguishable by $E_{z}$ just in case $\langle\Psi_{1}|\,E_{z}\,|\Psi_{1}\rangle\approx 0$ and $\langle\Psi_{2}|\,E_{z}\,|\Psi_{2}\rangle\approx 1$, or $\langle\Psi_{1}|\,E_{z}\,|\Psi_{1}\rangle\approx 1$ and $\langle\Psi_{2}|\,E_{z}\,|\Psi_{2}\rangle\approx 0$, i.e., the absolute difference in their probabilities is close to 1. The outcome $Z=z$ is a reliable indicator of the quantum state in the sense that, in nearly 100% of the worlds compatible with the outcome $Z=z$, it correctly indicates the actual quantum state. This suggests the following sufficient condition:
OI with respect to $E_{z}$
If the Born probability distributions of $E_{z}$ assigned by a set of quantum states are all within (say) $10^{-200}$ of each other, then observing the outcome $Z=z$ does not reliably distinguish among those states.
In what follows, unless specified otherwise, we shall understand OI with this criterion.
3.2 Observational Indistinguishability of Typical Quantum States
In this section, we show that distribution typicality implies observation typicality. We do this separately for OQM, BM, and GRW. The same arguments as for OQM and BM also provide OI for Everettian quantum mechanics.
(1) OI in OQM. Suppose we carry out any experiment $\mathscr{E}$; suppose $\mathscr{E}$ is associated, according to the Main Theorem about POVMs, with the POVM $E$; suppose we obtained the outcome $Z=z$ associated with the operator $E_{z}$. Then typical quantum states in a high-dimensional Hilbert space $\mathscr{H}_{0}$ (with $d_{0}>10^{1000}$) are not distinguished by this outcome from $\rho_{0}$.
Indeed, consider $\mathscr{H}_{0}$ with $d_{0}>10^{1000}$ and $\varepsilon=10^{-200}$. By Theorem 1 for any POVM element $E_{z}$, for ($1-\varepsilon)$-most $\Psi$ in $\mathscr{H}_{0}$, $|\langle\Psi|\,E_{z}\,|\Psi\rangle-\operatorname{tr}(\rho_{0}E_{z})|\leq\frac{1}{\sqrt{\varepsilon d_{0}}}<10^{-400}$. Hence, the Born probability distributions of $E_{z}$ assigned by ($1-\varepsilon)$-most $\Psi$ are all within $\tfrac{1}{2}\varepsilon$ of that of $\rho_{0}$, and thus within $\varepsilon$ of each other. As discussed above, a reliable distinction would require the observation of an event that has probability near 1 in one case and near 0 in the other, which does not happen here. Therefore, observing $E_{z}$ does not reliably distinguish those states from each other or from the normalized projection. So, we arrive at observation typicality in $\mathscr{H}_{0}$.
Again, we may think of $\mathscr{H}_{0}$ as the subspace demanded by the Past Hypothesis. Thus, for any observation $E_{z}$, typical quantum states in $\mathscr{H}_{0}$ are not distinguished by $E_{z}$ from the Wentaculus density matrix. Neither does it distinguish typical quantum states in $\mathscr{H}_{0}$ from each other.
In quantum theory with unitary evolution, this extends to later states too: Since typical $\Psi_{t}$ from the appropriately evolved subspace $U_{t}\mathscr{H}_{0}$ lead to nearly the same probability for outcome $Z=z$, the observation of this outcome does not distinguish among different $\Psi_{t}$.
(2) OI in BM. In BM, a key fact about OI is this:
the probability distribution of the configuration $Q_{t}$ of the universe at time $t$, i.e., the $|\Psi_{t}|^{2}$ distribution in configuration space, is nearly independent of $\Psi_{t_{0}}$ for typical $\Psi_{t_{0}}\in\mathbb{S}(\mathscr{H}_{0})$. (5)
This follows from Theorem 1 by taking the POVM $E$ to be the Heisenberg-evolved configuration observable, i.e., the projection-valued measure jointly diagonalizing all position operators at time $t$, or
$E(B)=U_{t}^{\dagger}1_{B}U_{t},$ (6)
where $B$ is any subset of configuration space and $1_{B}$ the multiplication operator multiplying by the characteristic function of $B$. Note that the probability distribution associated with this POVM is exactly the $|\Psi_{t}|^{2}$ distribution:

### v2

@@ 2.2 Main Motivation: The Past Hypothesis
In fact, it has been suggested for the explanation of the thermodynamic arrow of time (Feynman 1963, p. 115), (Penrose 1979, Lebowitz 1993, Albert 2000, Goldstein et al. 2020, Chen 2021) that the initial state of the universe must be restricted to a set of low-entropy states; in a quantum theory, such a set would be given by a suitable subspace $\mathscr{H}_{0}$ of the Hilbert space $\mathscr{H}$ of the universe, and the presumed additional law might then be formulated as follows:
@@
With respect to the normalized uniform measure $u_{0}$, we expect that typical wave functions in $\mathbb{S}(\mathscr{H}_{0})$ will evolve in a way such that $S_{qB}$ increases in the medium and the long run, satisfying the Developmental Conjecture formulated by Goldstein et al. (2020), a version of the Second Law of Thermodynamics for a quantum universe. In the philosophical literature, $\rho_{0}$ has been called the Wentaculus density matrix, corresponding to the initial state of the Wentaculus theory (Chen 2021, 2024a).
@@ 3.1 General Distinctions
There are various kinds of observational indistinguishability (OI). Here we clarify the distinction between in-principle OI and in-practice OI. Discussions in physics often focus on exact equivalence of predictions (or retrodictions) about observational outcomes. In general relativity, this is framed in terms of exact agreement (up to an isometry transformation) of the past lightcones of observers (Manchak 2009). In quantum foundations, many results pertain to exact equality of theoretical probability distributions. While exact equivalence is often sufficient for OI, it is not always necessary. Resource limitations mean that we sometimes reach in-practice OI long before hitting the in-principle limits. Nevertheless, sharp mathematical results are sometimes more easily proven by focusing on in-principle limits. Naturally, in-principle OI implies in-practice OI.
Observation typicality pertains to the stronger sense of OI—in principle, rather than merely in practice: No progress in technology would improve the situation, nor would more costly experiments or any other added resources. That is because we show that no substantial information about $\Psi_{t_{0}}$ can be obtained from any experiment using any resources, whenever the dimension $d_{0}$ of $\mathscr{H}_{0}$ is sufficiently large. It may seem surprising that an in-principle limitation can follow from an approximate equality (such as $\langle\Psi|E_{z}|\Psi\rangle\approx\operatorname{tr}(\rho_{0}E_{z})$), for two reasons. First, it would seem that nearby quantities can be distinguished better by measuring them with higher resolution; but here, the quantities are, not measurement outcomes, but probabilities of outcomes. Second, it would seem that probabilities can be measured more accurately if we repeat the experiment sufficiently often; but here, that is not possible, as we will discuss in §3.5.
3.2 Mathematical Descriptions
In §3.3, we will discuss for OQM, BM, GRW, and EQM which limitations for observers arise and how. But first, in this section, we describe the mathematical facts on which theses conclusions rest.
3.2.1 Not Reliably Distinguishable
We first show that, as a consequence of Theorem 1, typical vectors in $\mathscr{H}_{0}$ are not reliably distinguishable.
Let us make precise what we mean by “reliable.” Suppose we use outcome $z$ to determine whether $A$ is the case or $B$ is the case. We say that $E_{z}$$(1-\delta)$-reliably” distinguishes $A$ from $B$ just in case $|\mathbb{P}_{A}(Z=z)-\mathbb{P}_{B}(Z=z)|\geq 1-\delta$, where $\mathbb{P}_{A}$ means the probability if $A$ is the case, and likewise for $\mathbb{P}_{B}$. Thus, a reliability of 95% requires that the two probabilities differ by at least $95\%$. The reliability lies between 0 and 1, where 1 means perfectly reliable and 0 means completely unreliable. [^8] Under the decision rule “if $z$ occurs, guess $A$; otherwise guess $B$,” and assume $\mathbb{P}_{A}(z)\geq\mathbb{P}_{B}(z)$. Then $|\mathbb{P}_{A}(z)-\mathbb{P}_{B}(z)|\geq 1-\delta$ implies $\text{Sensitivity}=\mathbb{P}_{A}(z)\geq 1-\delta,$ and $\text{Specificity}=1-\mathbb{P}_{B}(z)\geq 1-\delta,$ so the Type I error (false-positive) rate $\mathbb{P}_{B}(z)\leq\delta$ and the Type II error (false-negative) rate $1-\mathbb{P}_{A}(z)\leq\delta$. Hence, smaller $\delta$ means smaller Type I and Type II errors in hypothesis testing.
Definition 1.
Let $\{E_{z}:z\in\mathscr{Z}\}$ be any POVM, $z$ any element of $\mathscr{Z}$, and $\delta>0$. We say $E_{z}$ $(1-\delta)$-reliably distinguishes the density matrices $\rho_{A},\rho_{B}$ if and only if $|\operatorname{tr}(\rho_{A}E_{z})-\operatorname{tr}(\rho_{B}E_{z})|\geq 1-\delta$.
Note that distinguishing at a reliability level $1-\delta$ also counts as distinguishing at any lower reliability level (i.e., larger $\delta$).
Example 3.
Consider again Example 2, with $E_{z}=P_{\uparrow}=\frac{1}{2}(I+\sigma_{z})$. If $\delta$ is taken to be $0.1,0.2,0.3$, then $E_{z}$ $(1-\delta)$-reliably distinguishes, respectively, $1\%,4\%,9\%$ of the total pairs of $\psi$’s in $\mathbb{C}^{2}$. (These fractions are given by $\delta^{2}$. [^9] Proof: Let $p_{A}=\langle\Psi_{A}|P_{\uparrow}|\Psi_{A}\rangle$ and likewise $p_{B}$. The set of $(p_{A},p_{B})\in[0,1]^{2}$ with $|p_{A}-p_{B}|\geq 1-\delta$ is the union of the triangles $(0,1-\delta)-(\delta,1)-(0,1)$ and $(1-\delta,0)-(1,\delta)-(1,0)$, which has area $\delta^{2}$. By Footnote 7, the set of $(\Psi_{A},\Psi_{B})$ with such $(p_{A},p_{B}$) has the same measure.)
We can show that for fixed $\delta$, in higher dimensional Hilbert spaces, Theorem 1 forces the fraction of pairs that $E_{z}$ $(1-\delta)$-reliably distinguishes to drop quickly and become essentially zero.
Corollary 1.
Let $\{E_{z}:z\in\mathscr{Z}\}$ be any POVM, $z$ any element of $\mathscr{Z}$, and $\varepsilon>0$. If $d_{0}>4/\varepsilon$, then for $(1-\varepsilon)$-most $\Psi_{A}\in\mathbb{S}(\mathscr{H}_{0})$, $E_{z}$ fails to $\frac{1}{2}$-reliably distinguish $\rho_{A}:=|\Psi_{A}\rangle\langle\Psi_{A}|$ from $\rho_{B}:=\rho_{0}$. Furthermore, If $d_{0}>16/\varepsilon$, then for $(1-\varepsilon)^{2}$-most choices of $\Psi_{A},\Psi_{B}$, $E_{z}$ fails to $\frac{1}{2}$-reliably distinguish $\rho_{A}$ from $\rho_{B}:=|\Psi_{B}\rangle\langle\Psi_{B}|$.
This follows from Theorem 1 via a few steps of algebra. [^10] Proof of the first statement: By (2), $|\operatorname{tr}(\rho_{A}E_{z})-\operatorname{tr}(\rho_{B}E_{z})|=|\langle\Psi_{A}|E_{z}|\Psi_{A}\rangle-\operatorname{tr}(\rho_{0}E_{z})|\leq\|E_{z}\|/\sqrt{\varepsilon d_{0}}=1/\sqrt{\varepsilon d_{0}}<1/2$. Second statement: $\Psi_{A}$ and $\Psi_{B}$ are chosen like independent random variables, so for $(1-\varepsilon)^{2}$-most pairs $(\Psi_{A},\Psi_{B})$, (2) for $B=E_{z}$ holds for both $\Psi$s, and thus $|\operatorname{tr}(\rho_{A}E_{z})-\operatorname{tr}(\rho_{B}E_{z})|=|\langle\Psi_{A}|E_{z}|\Psi_{A}\rangle-\langle\Psi_{B}|E_{z}|\Psi_{B}\rangle|\leq|\langle\Psi_{A}|E_{z}|\Psi_{A}\rangle-\operatorname{tr}(\rho_{0}E_{z})|+|\langle\Psi_{B}|E_{z}|\Psi_{B}\rangle-\operatorname{tr}(\rho_{0}E_{z})|\leq 2/\sqrt{\varepsilon d_{0}}<1/2$. Note that it trivially follows that $\rho_{A},\rho_{B}$ cannot be $(1-\delta)$-reliably distinguished for any $\delta$ between 0 and 1/2.
For example, for dimension $d_{0}=10^{100}$ and any $20<\delta<1/$ 2, $E_{z}$ can $(1-\delta)$-reliably distinguish at most the fraction $3.2\times 10^{-99}$ of the pairs of wave functions.
Corollary 1 should be contrasted with the statement that for large $d_{0}$ and purely randomly and independently chosen $\Psi_{A},\Psi_{B}\in\mathbb{S}(\mathscr{H}_{0})$, the two vectors will be nearly orthogonal, so there exists a POVM that distinguishes them reliably. Note, however, that if we keep this POVM $E$ fixed and consider new, random $\Psi_{A},\Psi_{B}$, then the two yield nearly the same Born distribution for $E$. So, the two statements differ in the order of quantifiers (“for most,” “there is”): for most $\Psi_{A},\Psi_{B}$, there is $E$ that reliably distinguishes $\Psi_{A},\Psi_{B}$; but according to Corollary 1, it is not the case that there is $E$ such that for most $\Psi_{A},\Psi_{B}$, $E$ reliably distinguishes $\Psi_{A},\Psi_{B}$.
Likewise, for every given $\Psi_{A}\in\mathbb{S}(\mathscr{H}_{0})$, there is a POVM $E$ that distinguishes $|\Psi_{A}\rangle\langle\Psi_{A}|$ $(1-d_{0}^{-1})$-reliably from $\rho_{0}$: just take $E_{A}=|\Psi_{A}\rangle\langle\Psi_{A}|$ and $E_{B}=I-E_{A}$. [^11] Proof: $\operatorname{tr}(\rho_{A}E_{A})=1$ and $\operatorname{tr}(\rho_{0}E_{B})=1-\operatorname{tr}(\rho_{0}E_{A})=1-\langle\Psi_{A}|\rho_{0}|\Psi_{A}\rangle=1-d_{0}^{-1}$.
Corollary 1 should also be contrasted with the problem of distinguishing between the case of a mixed state $\rho_{0}$ and the case of a random pure state $\Psi$ with distribution $u_{0}$; see §3.5.2 for details.
3.2.2 Not Even Unreliably Distinguishable
Next, we show that typical vectors in $\mathscr{H}_{0}$ are not even unreliably distinguishable if the observation isn’t too unlikely.
Consider an experiment with POVM $\{E_{z}:z\in\mathscr{Z}\}$ and outcome $Z\in\mathscr{Z}$, and consider two density matrices $\rho_{A},\rho_{B}$. Suppose that for a particular $z$,
$\mathbb{P}_{A}(Z=z)=\operatorname{tr}(\rho_{A}E_{z})=\frac{3}{4}~{}~{}~{}\text{and}~{}~{}~{}\mathbb{P}_{B}(Z=z)=\operatorname{tr}(\rho_{B}E_{z})=\frac{1}{4}.$ (8)
Suppose we observed the outcome $z$; then, although it does not distinguish reliably between $\rho_{A}$ and $\rho_{B}$ at a reliability level close to 1, one can argue that this outcome at least gives us some probabilistic information, as it favors option $A$ over option $B$ by a factor of 3. We now investigate how strongly our observations can favor $\Psi_{A}$.
Definition 2.
For any density matrices $\rho_{A},\rho_{B}$ on $\mathscr{H}$ and any POVM $\{E_{z}:z\in\mathscr{Z}\}$, we define the favoring ratio for $A$ given $z$ to be
$r_{A|z}:=\frac{\operatorname{tr}(\rho_{A}E_{z})}{\operatorname{tr}(\rho_{B}E_{z})}\,.$ (9)
The next statement expresses that in our case, the favoring ratio typically deviates from 1 by merely a negligible amount.
Corollary 2.
Let $\{E_{z}:z\in\mathscr{Z}\}$ be any POVM, $z$ any element of $\mathscr{Z}$, and $\varepsilon>0$. If $\operatorname{tr}(\rho_{0}E_{z})>\varepsilon$, then for $(1-\varepsilon)$-most $\Psi_{A}\in\mathbb{S}(\mathscr{H}_{0})$, the favoring ratio of $\rho_{A}=|\Psi_{A}\rangle\langle\Psi_{A}|$ over $\rho_{B}=\rho_{0}$ obeys
$1-\frac{1}{\sqrt{\varepsilon^{3}d_{0}}}\leq r_{A|z}\leq 1+\frac{1}{\sqrt{\varepsilon^{3}d_{0}}}\,.$ (10)
Furthermore, if also $d_{0}>9\varepsilon^{-3}$, then for $(1-\varepsilon)^{2}$-most choices of $\Psi_{A},\Psi_{B}$ in $\mathbb{S}(\mathscr{H}_{0})$, the favoring ratio of $\rho_{A}=|\Psi_{A}\rangle\langle\Psi_{A}|$ over $\rho_{B}=|\Psi_{B}\rangle\langle\Psi_{B}|$ obeys
$1-\frac{2}{\sqrt{\varepsilon^{3}d_{0}}}\leq r_{A|z}\leq 1+\frac{3}{\sqrt{\varepsilon^{3}d_{0}}}\,.$ (11)
In both cases, the favoring ratio is close to 1 whenever $d_{0}\gg\varepsilon^{-3}$.
This follows from Theorem 1 through some steps of algebra. [^12] Proof: By hypothesis, $\eta:=(\varepsilon d_{0})^{-1/2}/\operatorname{tr}(\rho_{0}E_{z})\leq 1/\sqrt{\varepsilon^{3}d_{0}}$. By Theorem 1, $(1-\varepsilon)$-most $\Psi_{A}$ obey (2) with $B=E_{z}$; dividing by $\operatorname{tr}(\rho_{0}E_{z})$, we obtain that $\langle\Psi_{A}|E_{z}|\Psi_{A}\rangle/\operatorname{tr}(\rho_{0}E_{z})\in[1-\eta,1+\eta]$. Second statement: By hypothesis, $\eta<1/3$. By Theorem 1, for $(1-\varepsilon)^{2}$-most pairs, both $\Psi$s obey (2) with $B=E_{z}$; we obtain the same relation as stated above for $\Psi_{A}$ also for $\Psi_{B}$. If $\langle\Psi_{A}|E_{z}|\Psi_{A}\rangle\geq\langle\Psi_{B}|E_{z}|\Psi_{B}\rangle$, then $1\leq\langle\Psi_{A}|E_{z}|\Psi_{A}\rangle/\langle\Psi_{B}|E_{z}|\Psi_{B}\rangle\leq(1+\eta)/(1-\eta)=1+2\eta/(1-\eta)<1+3\eta$ using $\eta<1/3$. If $\langle\Psi_{A}|E_{z}|\Psi_{A}\rangle\leq\langle\Psi_{B}|E_{z}|\Psi_{B}\rangle$, then $1\geq\langle\Psi_{A}|E_{z}|\Psi_{A}\rangle/\langle\Psi_{B}|E_{z}|\Psi_{B}\rangle\geq(1-\eta)/(1+\eta)=1-2\eta/(1+\eta)\geq 1-2\eta$. In words, it means that the outcome $z$, if it is not very improbable, leads to no significant probabilistic preference of $\Psi_{A}$ over $\rho_{0}$, or of $\rho_{0}$ over $\Psi_{A}$, or of $\Psi_{A}$ over $\Psi_{B}$, or of $\Psi_{B}$ over $\Psi_{A}$.
What about outcomes that are very improbable? They should not occur if (as we assume) the theory is correct, unless there are so many possible outcomes that the probabilities of the very improbable ones can add up to a significant value. That situation occurs in the example of 1000 coin tosses mentioned already near (6), so, unlike the situation considered by Borel, we need to also consider here events not specified in advance. That is, we need to reconsider the relevant size of $\varepsilon$ because $\varepsilon$ now appears in two roles: the maximal measure of the exceptional $\Psi$’s and the minimal probability of the improbable observations. So, how small could the probability of an observed outcome be? Like the coin toss example, an experiment can have a large number $L$ of possible outcomes, and if all have equal or similar probability, then these probabilities will be tiny of order $1/L$. If we need to distinguish between $L$ different observations, we can perhaps reach $L=10^{10^{10}}$ by collecting $10^{9}$ data points, each given by a number with 10 decimal digits. On the other hand, as of this writing (2025), the total storage capacity of all computers in the world is estimated at about $10^{26}$ bytes, so we would not be able to even record the outcome if $L$ exceeded $10^{10^{30}}$, which motivates us to consider,